The force of attraction F between two masses m and m separated by a distance g is given as F= (Gmm) ÷g×g where G is the universal gravitational constant. Use dimensions to derive S.I units of G
1
Expert's answer
2019-12-12T06:22:17-0500
"[F]=[G][m]^2[r]^{-2}"

"kg\cdot m\cdot s^{-2}=[G]kg^2\cdot m^{-2}"

Thus,


"[G]=kg^{-1}\cdot m^{3}\cdot s^{-2}"